The equation for a water line is y= -2/3x -12=
Under the -2 is a horizontal line to look like a fraction I couldn't put an underline under the -2.
Suppose you want to put a pink flamingo lawn ornament in your backyard, but you want to avoid placing it directly over the water line, in case you need to excavate the line for repairs in the future. Could you place it at the point
(-4,-10)?
Thankyou,
How would I solve this?

Answer by Alan3354(69443)

The equation for a water line is y= -2/3x -12=
Under the -2 is a horizontal line to look like a fraction I couldn't put an underline under the -2.
Suppose you want to put a pink flamingo lawn ornament in your backyard, but you want to avoid placing it directly over the water line, in case you need to excavate the line for repairs in the future. Could you place it at the point
(-4,-10)?
-------------------
y= -2/3x -12= (there's an = at the end??)
y= -2/3x -12 I'll use this.
I think this means there's a water line underground whose location is described by that equation.
The object is to see if the point (-4,-10) is above the water line.
Sub (-4) for x and see if y is -10.
y= -2/3x -12
y = (-2)*(-4)/3 - 12
y = 8/3 - 12
y = -28/3
That's 2/3 units more than -10, so it depends on the size of the ornament.